OpenStudy (anonymous):

Evaluate the following integrals by first reversing the order of integration. ∫(0,8)∫(y^(1/3),2) 8e^x^2 dxdy. Why is the limit when you reverse 0<=x<=2, 0<=y<=x^3 and not x^3<=y<=8?

OpenStudy (anonymous):

I'm confused with finding the limits for each when you reverse the order anyone got any tips?

OpenStudy (turingtest):

drawing out the region of integration is almost always a good idea

OpenStudy (turingtest):



OpenStudy (anonymous):

Yeah I got that too, but can't you say y^(1/3) to 8 is the limit? Go along the curve y=x^3 and stop at 8?

OpenStudy (turingtest):

well what would the area between y^(1/3) and y=8 look like?

OpenStudy (turingtest):

or rather x^3<=y<=8

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

it'd be the shaded area under y=8 but above y=x^3?

OpenStudy (anonymous):

which we're not looking for right

OpenStudy (turingtest):

you got it :)